OpenStudy (anonymous):

How do you simplify this ratio 9/27

OpenStudy (solomonzelman):

both the top and bottom are divisible by 9

OpenStudy (anonymous):

9/27 = 1/3. Divide numerator and denominator by 9, which is the greatest number that goes into both.

OpenStudy (sreekar369):

9/27=1/3=0.3333...........